# Unraveling the structure of $\Lambda$ hyperons with polarized $\Lambda\bar{\Lambda}$ pairs

### v1

@@ Abstract
With data collected in a dedicated off-resonance scan within the center-of-mass energy range of 2.3864--3.0800 GeV, the BESIII collaboration provides the first complete energy-dependent measurements of the $\Lambda$ electromagnetic form factors in the time-like region. By combining double-tag and single-tag events from the $e^{+}e^{-}\to\Lambda\bar{\Lambda}\to p\pi^{-}\bar{p}\pi^{+}$ reaction, we achieve a complete decomposition of the spin structure of the reaction at five energy points, with high statistical and systematic precision. Our data reveal that while the modulus of the ratio between the electric and magnetic form factor, $R(q^{2})=|G_{E}(q^{2})/G_{M}(q^{2})|$, remains fairly constant across the considered energy range, the relative phase $\Delta\Phi(q^{2})=\Phi_{E}(q^{2})-\Phi_{M}(q^{2})$ changes by more than 90o between 2.396 and 2.6544 GeV. Using a fit to our data based on dispersion relations, the complex form factor ratio is determined as a function of $q^{2}$ and the preferred solution has multiple zero-crossings in the complex plane. From the derivative of the ratio at $q^{2}=0$, the root-mean-squared charge radius of the $\Lambda$ is obtained. The two most probable solutions yield a negative root-mean-squared charge radius, indicating an asymmetric charge distribution where the $ds$ quark pair lies close to the center of the $\Lambda$ hyperon.
@@
To understand how almost massless quarks form massive hadrons is a challenging endeavor, involving the full complexity of the strong interaction and its emergent phenomena. Electromagnetic form factors (EMFFs) provide a powerful tool to study the strong interaction. Probed in virtual photon-hadron processes, EMFFs quantify hadron structure in terms of analytic functions of the squared four-momentum transfer $q^{2}$ carried by the photon. More than 70 years of detailed investigations employing elastic electron-proton scattering have unraveled the spatial distribution and motion of quarks and gluons inside the proton [1, 2]. In particular, the electric form factor constitutes an instrument to determine the charge radius, complementary to atomic spectroscopy methods [3]. Possible EMFF sign changes, or zero-crossings, at high $q^{2}$, are governed by the momentum dependence of the running strong coupling and mass functions of the quarks inside the hadron [4, 5]. Such zero-crossings are predicted for nucleons by Dyson-Schwinger calculations but are not yet experimentally established [6]. While neutron data elucidate the $u-d$ quark flavor dependence [7], the impact of the $\approx$20 times heavier strange $s$-quark remains rather uncharted territory. Addressing this question is challenging since the short lifetime (10-10 s) of strange hyperons make them unfeasible for elastic electron-hadron scattering experiments. However, their time-like EMFFs can be accessed through the process $e^{+}e^{-}\to\gamma^{*}\to Y\bar{Y}$, where $Y$ ($\bar{Y}$) denotes a hyperon (antihyperon). Whereas space-like ($q^{2}$ = $-Q^{2}<0$) EMFFs are real, the time-like ($q^{2}>0$) ones are complex with a phase reflecting the fact that the virtual photon can undergo quantum fluctuations into an intermediate multihadron state e.g. $2\pi$ or $3\pi$ [8].
@@
A dedicated off-resonance energy scan at center-of-mass system (CMS) energies between 2.0 and 3.08 GeV was performed in 2014-2015 at BESIII. From the resulting rich data, partial EMFF information has been obtained for protons [16], neutrons [17, 18], $\Sigma$ [19, 20, 21] and $\Xi$ [22, 23, 24]. Furthermore, exclusive double-tag analyses of the $e^{+}e^{-}\to\Lambda\bar{\Lambda}\to p\pi^{-}\bar{p}\pi^{+}$ and $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}\to p\pi^{0}\bar{p}\pi^{0}$ reactions, enabled the complete determination of the EMFF for $\Lambda$ at one energy [10] and for $\Sigma^{+}$ at two energies [25]. In this Letter, we combine the analysis of double-tag $e^{+}e^{-}\to\Lambda\bar{\Lambda}\to p\pi^{-}\bar{p}\pi^{+}$ with single-tag $e^{+}e^{-}\to\Lambda X\to p\pi^{-}X$ and $e^{+}e^{-}\to\bar{\Lambda}X\to\bar{p}\pi^{+}X$ at five different energies. This enables the complete determination of the $\Lambda$ EMFF at five energies within the range 2.3864-3.0800 GeV. In addition, we measure the production cross-section at nine energies. The integrated luminosities at all CMS energies are listed in Table 1.
@@
where $\mathcal{W}(R,\Delta\Phi,\theta,\theta_{p})$ is the joint decay distribution, $\tau=\frac{q^{2}}{4m^{2}_{\Lambda}}$ and $\alpha$ is the $\Lambda$ decay asymmetry parameter. The latter has been measured by BESIII [29, 30, 31, 32] and CLAS [33]. Due to the numerous new measurements, the Particle Data Group (PDG) [34] continuously updates the world average. For this work, we adopt the current average $\alpha\right>=(\alpha-\bar{\alpha})/2$ from the BESIII measurements [29, 30, 31, 32], $\alpha=0.754$, assuming CP symmetry i.e. $\alpha=-\bar{\alpha}$. This allows including measurements of $\bar{\alpha}$ for increased precision. The resulting value differs from the latest PDG value only on the sub-percent level, which is well beyond the precision of our phase measurement.
@@
In the event selection, the point of closest approach of charged tracks to the interaction point must be inside a cylindrical volume defined by the radial distance $V_{xy}\leq 10$ cm and the longitudinal distance $V_{z}\leq 30$ cm from the interaction point. The $z$-axis is the symmetry axis of the MDC, and the polar angle with respect to the $z$-axis must fulfill $|\cos\theta_{\text{lab}}|\leq 0.93$. Furthermore, we require $p\pi^{-}$ and $\bar{p}\pi^{+}$ combinations to have good vertex fits for production and decay, which determine the decay length $L$ as the distance between the two vertices. If more than one combination exists in an event, we select the one giving the largest $L$. This requirement selects the long-lived Λ/$\Lambda/\bar{\Lambda}$ while rejecting random or non-resonant $p\pi^{-}$ and $\bar{p}\pi^{+}$ combinations. The events are organized into three independent categories, referred to as double-tag $\Lambda\bar{\Lambda}$, single-tag $\Lambda$ and single-tag $\bar{\Lambda}$. The double-tag $\Lambda\bar{\Lambda}$ selection is essentially the same as in Ref. [10], but with complete particle identification (PID) of protons, antiprotons and pions, using ToF and $dE/dx$ information from the MDC. The $\chi^{2}_{4C}$ of a four-constraint (4C) kinematic fit, exploiting four-momentum conservation of the $e^{+}e^{-}\to p\pi^{-}\bar{p}\pi^{+}$ process, must be less than 130. Finally, the $p\pi^{-}$ and $\bar{p}\pi^{+}$ combinations must satisfy $|M(p\pi)-m_{\Lambda}|\leq 6$ MeV/$c^{2}$ which corresponds to ±$\pm 4\sigma$ of the peak in the $p\pi$ invariant mass spectrum. These criteria give efficiencies within the range of $(13-22)\%$ (see $\epsilon_{\Lambda\bar{\Lambda}}$ in Table 1) and a background fraction of less than $1\%$.
@@
For single-tag events, we do not exploit information from the untagged Λ/$\Lambda/\bar{\Lambda}$, decaying either via the charged mode or through the neutral $\Lambda\to n\pi^{0}/\bar{\Lambda}\to\bar{n}\pi^{0}$. The efficiencies for the single-tag modes, denoted $\epsilon_{1,\Lambda}$, $\epsilon_{2,\Lambda}$, $\epsilon_{1,\bar{\Lambda}}$ and $\epsilon_{2,\bar{\Lambda}}$, where 1 refers to the charged mode and 2 to the neutral one, are listed in Table 1. We note that the efficiencies $\epsilon_{2,\Lambda}$ and $\epsilon_{2,\bar{\Lambda}}$ are larger since there is no overlap with the double-tag sample. The background, estimated using the $M(p\pi^{-})$ and $M(\bar{p}\pi^{+})$ sidebands (see [37]), is found to be $\leq$1.5and $\leq$0.5for single-tag $\Lambda$ and single-tag $\bar{\Lambda}$, respectively. The small background in both double-tag and single-tag data samples is confirmed in simulations of the inclusive $e^{+}e^{-}\to q\bar{q}$ process and reactions with similar topology, e.g. $e^{+}e^{-}\to\Lambda\bar{\Sigma}^{0}+c.c.$.
Table 1: CMS energies (in GeV), integrated luminosities $\mathcal{L}$ (in pb-1), ISR+VP corrections $(1+\delta)$, background subtracted yields $N$, efficiencies $\epsilon$ (in %) and the resulting Born cross-sections (in pb) with statistical and systematic uncertainties. For the single-tag efficiencies $\epsilon_{1,\Lambda}$, $\epsilon_{2,\Lambda}$, $\epsilon_{1,\bar{\Lambda}}$ and $\epsilon_{2,\bar{\Lambda}}$, 1 and 2 refer to the charged and neutral decay modes, respectively, of the untagged decay.
Energy $\mathcal{L}$ $(1+\delta)$ $N_{\Lambda\bar{\Lambda}}$ $N_{\bar{\Lambda}}$ $N_{\Lambda}$ $\epsilon_{\Lambda\bar{\Lambda}}$ $\epsilon_{1,\bar{\Lambda}}$ $\epsilon_{2,\bar{\Lambda}}$ $\epsilon_{1,\Lambda}$ $\epsilon_{2,\Lambda}$ $\sigma_{Born}$ 2.3864 22.6 0.98 $227\pm 15$ $373\pm 19$ $315\pm 20$ $18.6$ $12.3$ $31.8$ $11.7$ 28.2 $132.7(45)(28)$ 2.3960 66.9 0.99 $596\pm 24$ $973\pm 32$ $575\pm 25$ 18.5 12.1 31.4 7.5 18.4 $120.2(26)(25)$ 2.6544 67.8 1.23 $224\pm 15$ $300\pm 18$ $244\pm 17$ 21.7 9.7 31.2 9.0 27.7 $31.0(11)(07)$ 2.9000 105.5 1.57 $117\pm 11$ $120\pm 11$ $108\pm 11$ 19.1 6.5 23.5 6.2 22.0 $8.7(05)(03)$ 2.9500 16.0 1.65 $16\pm 4$ $12\pm 4$ $14\pm 4$ 17.4 6.4 23.8 6.1 21.9 $7.2(11)(02)$ 2.9810 16.0 1.71 $16\pm 4$ $13\pm 4$ $13\pm 4$ 18.8 5.9 22.9 5.6 21.5 $7.7(12)(02)$ 3.0000 15.8 1.75 $13\pm 3$ $12\pm 4$ $14\pm 4$ 12.9 6.2 23.2 5.9 21.6 $7.1(12)(02)$ 3.0200 17.3 1.75 $11\pm 3$ $12\pm 4$ $14\pm 4$ 14.1 6.1 23.8 5.8 22.2 $5.6(10)(02)$ 3.0800 126.6 1.85 $65\pm 8$ $60\pm 8$ $62\pm 8$ 14.8 4.1 16.1 3.9 15.1 $4.6(04)(02)$
@@
$\sigma_{B}=\frac{N_{\Lambda\bar{\Lambda}}}{\mathcal{L}(1+\delta)\epsilon_{\Lambda\bar{\Lambda}}\mathcal{B}_{1}^{2}},$ (2)
@@
$\sigma_{B}=\frac{N_{\Lambda}}{\mathcal{L}(1+\delta)\mathcal{B}_{1}\left(\epsilon_{1,\Lambda}\mathcal{B}_{1}+\epsilon_{2,\Lambda}\mathcal{B}_{2}\right)},$ (3)
@@
The ISR and VP correction factors are obtained from ConExc in an iterative procedure where the input lineshape of the initial step is obtained by applying a pQCD model [39] to data from BaBar [15] and BESIII [40]. Applying the resulting correction factors to our data, we perform a new fit using a dipole model [28]. The resulting refined lineshape is used as input in the next step. This procedure is repeated until the difference between the last two iterations becomes negligible compared to other uncertainties; typically (0.1-0.6)%. The cross-sections obtained for the three independent samples (double-tag $\Lambda\bar{\Lambda}$, single-tag $\Lambda$ and single-tag $\bar{\Lambda}$) all agree within the statistical uncertainty and are combined using the weighted mean. The results are listed in Table 1.
General systematic effects arise from the luminosity (1% at each energy point [41]) and $\mathcal{B}_{1,2}$ [34]. Systematic uncertainties specific to this analysis are evaluated for the relatively large samples at 2.396 GeV and 2.9 GeV [28], while obtained by linear interpolation or extrapolation in the remaining data points. Our analysis reveals small but non-negligible effects from differences at the sub-percent level in the result when varying the $\chi^{2}_{\rm{4C}}$ (double-tag) and $\Lambda$/$\bar{\Lambda}$ momentum (single-tag) selection requirements. In addition, there is a small uncertainty of (0.1-0.8)in the $\Lambda$/$\bar{\Lambda}$ reconstruction efficiency and above 2.9 GeV, a contribution of (0.5-1.2)from the $\Lambda\bar{\Sigma}^{0}$ background. Since the MC data used for the efficiency correction are generated with an angular distribution governed by $R$, the uncertainty in the input parameter $R$ will propagate to the cross-section. For energies above 2.9 GeV, this uncertainty is fairly large i.e. (2.2-2.8)%, while only within (0.6-1.4)below 2.9 GeV. The total systematic uncertainty is well below the statistical uncertainty in all data points except at 2.396 GeV, where it is of similar magnitude, see Table 1.
@@
$\begin{split}|G_{\text{eff}}(q^{2})|\equiv&\sqrt{\frac{\sigma_{\Lambda\bar{\Lambda}}(q^{2})}{(1+\frac{1}{2\tau})(\frac{4\pi\alpha^{2}_{\text{EM}}\beta}{3q^{2}})}}.\end{split}$ (4)
@@
Figure 1: The effective form factor $G_{\text{eff}}$ from this work (black dots), from BESIII ISR [42], BESIII energy scan [43, 10] and BaBar ISR [15]. The horizontal error bars denote the bin size.
@@
Figure 2: Top: The modulus $R=|G_{E}/G_{M}|$ from this work (black dots), from a previous BESIII study (white dot) [10] and from BaBar (gray dots) [15]. The band represents a theoretical analysis outlined in Ref. [8], showing the most probable scenario $[N_{\text{th}},N_{\text{asy}}]=[0,3]$. The BESIII 2019 data point is superseded by the point at the same CMS energy in this work and is therefore not included in the fit. The red dashed line marks the theoretical threshold $q^{2}_{th}$ and the red dashed-dotted line the physical threshold $q^{2}_{\text{phys}}$. Bottom: The corresponding measurements of the relative phase $\Delta\Phi$.
@@
Systematic effects are investigated by varying all selection criteria for all samples. In most cases the effects are negligible, except for the $M(p\pi)$ window that slightly influences $\Delta\Phi$, and the requirement on the $\chi^{2}$ from the vertex fit that has a small impact on $R$. In addition, there is an uncertainty in the $\Lambda$/$\bar{\Lambda}$ reconstruction efficiency for both $R$ and $\Delta\Phi$. Background effects are studied by sampling MC data mimicking the sideband events in real data. These background samples are added to signal MC events in 1000 MLL fits. We find the effect to be small, though slightly larger for $R$ than for $\Delta\Phi$. The systematic uncertainties of $R$ and $\Delta\Phi$ are summarized in Table 3.
@@
Table 3: Systematic uncertainties in $R$ ($\Delta\Phi$) in % at each CMS energy (GeV). The sources are from (I) the $M(p\pi)$ window, (II) vertex fit $\chi^{2}$ requirement, (III) Λ/$\Lambda/\bar{\Lambda}$ reconstruction efficiency, and (IV) background effects.
@@
where $q^{2}_{\text{th}}$ corresponds to the squared-mass of the lowest-lying hadronic state that can couple to the virtual photon and the $\Lambda\bar{\Lambda}$ final state, i.e. $\pi^{+}\pi^{-}\pi^{0}$, and $q^{2}_{\text{asy}}$ is the scale where the phase $\Delta\Phi(q^{2})$ approaches an integer multiple of $\pi$. At the physical threshold $q^{2}_{\text{phys}}=4m_{\Lambda}^{2}$, only $s$-waves can contribute to the $\Lambda\bar{\Lambda}$ production and therefore $R(q^{2}_{\text{phys}})\overset{!}{=}1$ and $\Delta\Phi(q^{2}_{\text{phys}})\overset{!}{=}0$. The number of zero-crossings of the complex ratio $G_{E}(q^{2})/G_{M}(q^{2})$ is given by $\Delta N=N_{\text{asy}}-N_{\text{th}}$ but their location in the complex plane cannot be determined from this approach. In the calculations, the imaginary part of the ratio is parameterized by Chebyshev polynomials of the first kind. Seven or eight degrees appear optimal and give consistent results as shown in Table 4. The most probable scenario, $[N_{\text{th}},N_{\text{asy}}]=[0,3]$ ($P=85\%)$ shown in Fig. 2, is strikingly more probable than the second scenario $[N_{\text{th}},N_{\text{asy}}]=[0,4]$. Notably, all possible solutions imply multiple zero-crossings of $G_{E}(q^{2})/G_{M}(q^{2})$, in line with Dyson-Schwinger-based predictions for nucleons [4] and vector mesons [45].
@@
For neutral baryons with an anomalous magnetic moment, $\langle r^{\Lambda}_{E}\rangle^{2}$ can be calculated from the derivative of $R(q^{2})$ at $q^{2}=0$ [8]. The results are presented in Table 4, including the probability of each solution as calculated using Monte Carlo simulations [8]. The most probable $[N_{\text{th}},N_{\text{asy}}]$ scenario indicates a negative RMS charge radius ($\bar{r}_{E}^{\Lambda}=-0.076\pm 0.043$ fm for a seven degree polynomial). The effect is $\approx$2 standard deviations and the conclusion is valid also when lower degree polynomials are used in the fit. This indicates that the negative $sd$ quark pair in the $\Lambda$ is close to the center of the $\Lambda$, similar to the $dd$ pair in the neutron [46]. In summary, our measurements of time-like EMFFs, combined with dispersive relations, open new avenues to study the properties of unstable matter containing strange quarks and beyond.

### v2

@@ Abstract
With data collected in a dedicated energy scan from 2.3864 up to 3.0800 GeV, the BESIII collaboration provides the first complete energy-dependent measurements of the $\Lambda$ electromagnetic form factors in the time-like region. By combining double-tag and single-tag events from the $e^{+}e^{-}\to\Lambda\bar{\Lambda}\to p\pi^{-}\bar{p}\pi^{+}$ reaction, we achieve a complete decomposition of the spin structure of the reaction at five energy points, with high statistical and systematic precision. Our data reveal that while the modulus of the ratio between the electric and magnetic form factor, $R(q^{2})=|G_{E}(q^{2})/G_{M}(q^{2})|$, remains fairly constant across the considered energy range, the relative phase $\Delta\Phi(q^{2})=\Phi_{E}(q^{2})-\Phi_{M}(q^{2})$ changes by more than 90o between 2.396 and 2.6544 GeV. Using a fit to our data based on dispersion relations, the complex form factor ratio is determined as a function of $q^{2}$ and the preferred solution has multiple zero-crossings in the complex plane. From the derivative of the ratio at $q^{2}=0$, the root-mean-squared charge radius of the $\Lambda$ is obtained. The two most probable solutions yield a negative root-mean-squared charge radius, indicating an asymmetric charge distribution where the $ds$ quark pair lies close to the center of the $\Lambda$ hyperon.
@@
To understand how almost massless quarks form massive hadrons is a challenging endeavor, involving the full complexity of the strong interaction and its emergent phenomena. Electromagnetic form factors (EMFFs) provide a powerful tool to study the strong interaction. Probed in virtual photon-hadron processes, EMFFs quantify hadron structure in terms of analytic functions of the squared four-momentum transfer $q^{2}$ carried by the photon. More than 70 years of detailed investigations employing elastic electron-proton scattering have unraveled the spatial distribution and motion of quarks and gluons inside the proton [40, 36]. In particular, the electric form factor constitutes an instrument to determine the charge radius, complementary to atomic spectroscopy methods [29]. Possible EMFF sign changes, or zero-crossings, at high $q^{2}$, are governed by the momentum dependence of the running strong coupling and mass functions of the quarks inside the hadron [43, 42]. Such zero-crossings are predicted for nucleons by Dyson-Schwinger calculations but are not yet experimentally established [39]. While neutron data elucidate the udu-d quark flavor dependence [26], the impact of the $\approx$20 times heavier strange $s$-quark remains rather uncharted territory. Addressing this question is challenging since the short lifetime (10-10 s) of strange hyperons make them virtually unfeasible for elastic electron-hadron scattering experiments - the only existing measurement of this kind is for the $\Sigma^{-}$ hyperon [30]. However, time-like EMFFs of hyperons can be accessed through the process $e^{+}e^{-}\to\gamma^{*}\to Y\bar{Y}$, where $Y$ ($\bar{Y}$) denotes a hyperon (antihyperon). Whereas space-like ($q^{2}$ = $Q^{2}<0$) EMFFs are real, the time-like ($q^{2}>0$) ones are complex with a phase reflecting the fact that the virtual photon can undergo quantum fluctuations into an intermediate multihadron state e.g. $2\pi$ or $3\pi$ [35].
@@
A dedicated energy scan at center-of-mass system (CMS) energies between 2.0 and 3.08 GeV was performed in 2014-2015 at BESIII. At these energies, we expect the influence from vector charmonium resonances to be negligible and the production mechanism to be dominated by one-photon exchange. Each energy corresponds to a $q^{2}$ through the relation $E_{\text{CMS}/}^{2}/c^{2}=q^{2}$. From the resulting rich data, partial EMFF information has been obtained for protons [13], neutrons [18, 20], $\Sigma$ [17, 7, 11] and $\Xi$ [14, 16, 15]. Furthermore, exclusive double-tag analyses of the $e^{+}e^{-}\to\Lambda\bar{\Lambda}\to p\pi^{-}\bar{p}\pi^{+}$ and $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}\to p\pi^{0}\bar{p}\pi^{0}$ reactions, enabled the complete determination of the EMFF for $\Lambda$ at one energy [4] and for $\Sigma^{+}$ at two energies [10]. In this Letter, we combine the analysis of double-tag $e^{+}e^{-}\to\Lambda\bar{\Lambda}\to p\pi^{-}\bar{p}\pi^{+}$ with single-tag $e^{+}e^{-}\to\Lambda X\to p\pi^{-}X$ and $e^{+}e^{-}\to\bar{\Lambda}X\to\bar{p}\pi^{+}X$ at five different energies. This enables the complete determination of the $\Lambda$ EMFF at five energies within the range 2.3864-3.0800 GeV. In addition, we measure the production cross-section at nine energies. The integrated luminosities at all CMS energies are listed in Table 1.
@@
where $\mathcal{W}(R,\Delta\Phi,\theta,\theta_{p})$ is the joint decay distribution, $\tau=\frac{q^{2}}{4m^{2}_{\Lambda}}$ and $\alpha$ is the $\Lambda$ decay asymmetry parameter. The latter has been measured by BESIII [5, 19, 8, 21] and CLAS [32]. Due to the numerous new measurements, the Particle Data Group (PDG) [45] continuously updates the world average. For this work, we adopt the current average $\left<\alpha\right>=(\alpha-\bar{\alpha})/2$ from the BESIII measurements [5, 19, 8, 21], $\alpha=0.754$, assuming CP symmetry i.e. $\alpha=-\bar{\alpha}$. This allows including measurements of $\bar{\alpha}$ for increased precision. The resulting value differs from the latest PDG value only on the sub-percent level, which is well beyond the precision of our phase measurement.
@@
[…]
For single-tag events, we do not exploit information from the untagged $\Lambda/\bar{\Lambda}$, decaying either via the charged mode or through the neutral $\Lambda\to n\pi^{0}/\bar{\Lambda}\to\bar{n}\pi^{0}$. The efficiencies for the single-tag modes, denoted $\epsilon_{1,\Lambda}$, $\epsilon_{2,\Lambda}$, $\epsilon_{1,\bar{\Lambda}}$ and $\epsilon_{2,\bar{\Lambda}}$, where 1 refers to the charged mode and 2 to the neutral one, are listed in Table 1. We note that the efficiencies $\epsilon_{2,\Lambda}$ and $\epsilon_{2,\bar{\Lambda}}$ are larger since there is no overlap with the double-tag sample. The background, estimated using the $M(p\pi^{-})$ and $M(\bar{p}\pi^{+})$ sidebands (see [22]), is found to be $\leq$1.5% and $\leq$0.5% for single-tag $\Lambda$ and single-tag $\bar{\Lambda}$, respectively. The small background in both double-tag and single-tag data samples is confirmed in simulations of the inclusive $e^{+}e^{-}\to q\bar{q}$ process and reactions with similar topology, e.g. $e^{+}e^{-}\to\Lambda\bar{\Sigma}^{0}+c.c.$.
Table 1: CMS energies (in GeV), integrated luminosities $\mathcal{L}$ (in pb-1), ISR+VP corrections $(1+\delta)$, background subtracted yields NN, efficiencies $\epsilon$ (in %) and the resulting Born cross-sections (in pb) with statistical and systematic uncertainties. For the single-tag efficiencies $\epsilon_{1,\Lambda}$, $\epsilon_{2,\Lambda}$, $\epsilon_{1,\bar{\Lambda}}$ and $\epsilon_{2,\bar{\Lambda}}$, 1 and 2 refer to the charged and neutral decay modes, respectively, of the untagged decay.
Energy $\mathcal{L}$ $(1+\delta)$ $N_{\Lambda\bar{\Lambda}}$ $N_{\bar{\Lambda}}$ $N_{\Lambda}$ $\epsilon_{\Lambda\bar{\Lambda}}$ $\epsilon_{1,\bar{\Lambda}}$ $\epsilon_{2,\bar{\Lambda}}$ $\epsilon_{1,\Lambda}$ $\epsilon_{2,\Lambda}$ $\sigma_{B}$ 2.3864 22.6 0.98 $227\pm 15$ $373\pm 19$ $315\pm 20$ $18.6$ $12.3$ $31.8$ $11.7$ 28.2 $132.7(45)(28)$ 2.3960 66.9 0.99 $596\pm 24$ $973\pm 32$ $575\pm 25$ 18.5 12.1 31.4 7.5 18.4 $120.2(26)(25)$ 2.6544 67.8 1.23 $224\pm 15$ $300\pm 18$ $244\pm 17$ 21.7 9.7 31.2 9.0 27.7 $31.0(11)(07)$ 2.9000 105.5 1.57 $117\pm 11$ $120\pm 11$ $108\pm 11$ 19.1 6.5 23.5 6.2 22.0 $8.7(05)(03)$ 2.9500 16.0 1.65 $16\pm 4$ $12\pm 4$ $14\pm 4$ 17.4 6.4 23.8 6.1 21.9 $7.2(11)(02)$ 2.9810 16.0 1.71 $16\pm 4$ $13\pm 4$ $13\pm 4$ 18.8 5.9 22.9 5.6 21.5 $7.7(12)(02)$ 3.0000 15.8 1.75 $13\pm 3$ $12\pm 4$ $14\pm 4$ 12.9 6.2 23.2 5.9 21.6 $7.1(12)(02)$ 3.0200 17.3 1.75 $11\pm 3$ $12\pm 4$ $14\pm 4$ 14.1 6.1 23.8 5.8 22.2 $5.6(10)(02)$ 3.0800 126.6 1.85 $65\pm 8$ $60\pm 8$ $62\pm 8$ 14.8 4.1 16.1 3.9 15.1 $4.6(04)(02)$
@@
$\sigma_{B}=\frac{\sigma_{exp}}{1+\delta}=\frac{N_{\Lambda\bar{\Lambda}}}{\mathcal{L}(1+\delta)\epsilon_{\Lambda\bar{\Lambda}}\mathcal{B}_{1}^{2}},$ (2)
@@
$\sigma_{B}=\frac{\sigma_{exp}}{1+\delta}=\frac{N_{\Lambda}}{\mathcal{L}(1+\delta)\mathcal{B}_{1}\left(\epsilon_{1,\Lambda}\mathcal{B}_{1}+\epsilon_{2,\Lambda}\mathcal{B}_{2}\right)},$ (3)
@@
The ISR and VP correction factors are obtained from ConExc in an iterative procedure where the input lineshape of the initial step is obtained by applying a pQCD model [37] to data from BaBar [24] and BESIII [3]. Applying the resulting correction factors to our data, we perform a new fit using a dipole model [44]. The resulting refined lineshape is used as input in the next step. This procedure is repeated until the difference between the last two iterations becomes negligible compared to other uncertainties. The cross-sections obtained for the three independent samples (double-tag $\Lambda\bar{\Lambda}$, single-tag $\Lambda$ and single-tag $\bar{\Lambda}$) all agree within the statistical uncertainty and are combined using the weighted mean. The results are listed in Table 1.
General systematic effects arise from the luminosity (1% at each energy point [2]) and $\mathcal{B}_{1,2}$ [45]. Systematic uncertainties specific to this analysis are evaluated for the relatively large samples at 2.396 GeV and 2.9 GeV [44], while obtained by linear interpolation or extrapolation in the remaining data points. Our analysis reveals small but non-negligible effects from differences at the sub-percent level in the result when varying the $\chi^{2}_{\rm{4C}}$ (double-tag) and $\Lambda$$\bar{\Lambda}$ momentum (single-tag) selection requirements. In addition, there is a small uncertainty of (0.1-0.8)% in the $\Lambda$$\bar{\Lambda}$ reconstruction efficiency and above 2.9 GeV, a contribution of (0.5-1.2)% from the $\Lambda\bar{\Sigma}^{0}$ background. The uncertainty from the ISR and VP corrections were evaluated by using a pQCD fit for the cross-section lineshape. The difference between the dipole and pQCD model results, (0.1-0.6)%, is quoted as the uncertainty. Since the MC data used for the efficiency correction are generated with an angular distribution governed by $R$, the uncertainty in the input parameter $R$ will propagate to the cross-section. For energies above 2.9 GeV, this uncertainty is fairly large i.e. (2.2-2.8)%, while only within (0.6-1.4)% below 2.9 GeV. The total systematic uncertainty is well below the statistical uncertainty in all data points except at 2.396 GeV, where it is of similar magnitude, see Table 1.
@@
$\begin{split}|G_{\text{eff}}(q^{2})|\equiv&\sqrt{\frac{\sigma_{B}(q^{2})}{(1+\frac{1}{2\tau})(\frac{4\pi\alpha^{2}_{\text{EM}}\beta}{3q^{2}})}}.\end{split}$ (4)
@@
Figure 1: The effective form factor $G_{\text{eff}}$ from this work (black dots), BESIII ISR [9], BESIII energy scan [12, 4] and BaBar ISR [24]. The horizontal error bars denote the bin size.
@@
Figure 2: Top: The modulus $R=|G_{E}/G_{M}|$ from this work (black dots), from a previous BESIII study (white dot) [4] and from BaBar (gray dots) [24]. The band represents a theoretical analysis outlined in Ref. [35], showing the most probable scenario $N_{\text{th}},N_{\text{asy}}]=[0,3]$. The BESIII 2019 data point is superseded by the point at the same CMS energy in this work and is therefore not included in the fit. The red dashed line marks the theoretical threshold $q^{2}_{th}$ and the red dashed-dotted line the physical threshold $q^{2}_{\text{phys}}$. Bottom: The corresponding measurements of the relative phase $\Delta\Phi$.
@@
Systematic effects are investigated by varying all selection criteria for all samples. In most cases the effects are negligible, except for the $M(p\pi)$ window that slightly influences $\Delta\Phi$, and the requirement on the $\chi^{2}$ from the vertex fit that has a small impact on $R$. In addition, there is an uncertainty in the $\Lambda$$\bar{\Lambda}$ reconstruction efficiency for both $R$ and $\Delta\Phi$. Background effects are studied by sampling MC data mimicking the sideband events in real data. These background samples are added to signal MC events in 1000 MLL fits. We find the effect to be small, though slightly larger for $R$ than for $\Delta\Phi$. The systematic uncertainties of $R$ and $\Delta\Phi$ are summarized in Table 3.
@@
Table 3: Systematic uncertainties in $R$ ($\Delta\Phi$) in % at each CMS energy (GeV). The sources are from (I) the $M(p\pi)$ window, (II) vertex fit $\chi^{2}$ requirement, (III) $\Lambda/\bar{\Lambda}$ reconstruction efficiency, and (IV) background effects.
@@
where $q^{2}_{\text{th}}$ corresponds to the squared-mass of the lowest-lying hadronic state that can couple to the virtual photon and the $\Lambda\bar{\Lambda}$ final state, i.e. $\pi^{+}\pi^{-}\pi^{0}$, and $q^{2}_{\text{asy}}$ is the scale where the phase $\Delta\Phi(q^{2})$ approaches an integer multiple of $\pi$. At the physical threshold $q^{2}_{\text{phys}}=4m_{\Lambda}^{2}$, only $s$-waves can contribute to the $\Lambda\bar{\Lambda}$ production and therefore $R(q^{2}_{\text{phys}})\overset{!}{=}1$ and $\Delta\Phi(q^{2}_{\text{phys}})\overset{!}{=}2k\pi$. The number of zero-crossings of the complex ratio $G_{E}(q^{2})/G_{M}(q^{2})$ is given by $\Delta N=N_{\text{asy}}-N_{\text{th}}$, following Cauchy’s argument principle. In the calculations, the imaginary part of the ratio is parameterized by Chebyshev polynomials of the first kind. Seven or eight degrees appear optimal and give consistent results as shown in Table 4. The $\chi^{2}/\nu$ is obtained from the best fit result. To calculate the probability, we perform simulations of pseudo-experiments of different $N_{\text{th}},N_{\text{asy}}]$ scenarios, using the measured values and their uncertainties as input. The probability $P$ is given by the relative occurrence of a $N_{\text{th}},N_{\text{asy}}]$ scenario for a large number of pseudo-experiments. The most probable scenario, $N_{\text{th}},N_{\text{asy}}]=[0,3]$ ($P=85\%)$ shown in Fig. 2, is strikingly more probable than the second scenario $N_{\text{th}},N_{\text{asy}}]=[0,4]$. Notably, all possible solutions imply multiple zero-crossings of $G_{E}(q^{2})/G_{M}(q^{2})$, in line with Dyson-Schwinger-based predictions for nucleons [43] and vector mesons [46].
@@
For neutral baryons with an anomalous magnetic moment, $\langle r^{\Lambda}_{E}\rangle^{2}$ can be calculated from the derivative of $R(q^{2})$ at $q^{2}=0$ [35]. The results are presented in Table 4, including the probability of each solution as calculated using Monte Carlo simulations [35]. The most probable $N_{\text{th}},N_{\text{asy}}]$ scenario indicates a negative RMS charge radius ($\bar{r}_{E}^{\Lambda}=-0.076\pm 0.043$ fm for a seven degree polynomial). The effect is $\approx$2 standard deviations and the conclusion is valid also when lower degree polynomials are used in the fit. This indicates that the negative $sd$ quark pair in the $\Lambda$ is close to the center of the $\Lambda$, similar to the $dd$ pair in the neutron [23]. In summary, our measurements of time-like EMFFs, combined with dispersive relations, open new avenues to study the properties of unstable matter containing strange quarks and beyond.